Chapter 3, Section 3.2, Question 029 Using the vertex form, find a formula for the parabola with vertex (7,9) that passes thr I really need some help with this question please amd if you could explain how you got the answer. thank you!
Answer #1

Let parabola equation in Quadratse fossy and bxtc Yz (azo verten (7,9 nie -b za = 7 -b2 14 a 14a+bzo 0 and 9= a(72+ 7676 499+Then -b= 14(-7) ba १४ From @ C = -334 Then yz - 7x² +988 -334 =-7(x²_142 +49-49)-334 - 7(x-7)% +9. (y-9) 2-7 (x-7) 2 Cascanne